OpenStudy (laura.swart):

f(x)= (x-1)/(x^2+2x-8)

OpenStudy (laura.swart):

\[\frac{ x-1 }{x^2+2x-8}\]

OpenStudy (3mar):

May I help?

Nnesha (nnesha):

what's the question ??

jhonyy9 (jhonyy9):

like a first step x^2 +2x -8 = (x +4)(x-2) but what is the question of this exercise ?

OpenStudy (3mar):

Is the question is: simplify or find the root or determine the domain and the range of this function? Pick one!

Nnesha (nnesha):

or maybe find horizontal vertical asymptotes...